OpenStudy (samigupta8):

Let O,P,Q,R,S (O being origin) be five points all lying on a circle in order. P,Q,R,S have respective position vector p,q,r,s

OpenStudy (samigupta8):

∆POQ.∆ROS + ∆POS.QOR =? A).∆PRQ ∆PRS B).∆POR ∆QOS C).([]OPQR)([]QQRS) D).([]PQRS)^2 {[] denotes AREA of quadrilateral AND ∆denotes area of triangle}

OpenStudy (samigupta8):

I m just not getting this .Can we prove this part? (p*q)•(r*s)+(p*s)•(q*r)=(p*r)•(q*s)

Parth (parthkohli):

Assume that \(P_i = \left(r\cos \theta_i, r \sin \theta_i \right)\) and then check.

OpenStudy (samigupta8):

Really! That way.

Parth (parthkohli):

Yeah. That's a vector by the way.

OpenStudy (samigupta8):

Yep,that's a vector and do you intend to say that i should solve for each one of them individually (like taking dot product of each of them).

Parth (parthkohli):

Unfortunately...

OpenStudy (samigupta8):

I think that it is really very long way to go with. Didn't try that(obviously). I thought some STP would work out here in better way.

Parth (parthkohli):

I haven't studied vector geometry or stuff. Here's another way: PQRS is a cyclic quadrilateral. What do we know about cyclic quadrilaterals? :D

OpenStudy (samigupta8):

Not much ..(the geometry part is really very boring to me) Just this,the opposite angles are supplementary.

Parth (parthkohli):

Yeah. I don't know... check if that works, maybe?

OpenStudy (samigupta8):

@ganeshie8